OpenStudy (anonymous):

* Advance Algebra Financial Application * Geoffrey needs to purchase baseballs and bats for his youth baseball league. He’ll need at least 30 pieces of equipment, but he only has $870 to spend. Baseballs cost $20 each and bats cost $35 each. Which is a combination of baseballs and bats that Geoffrey can afford to buy? 14 baseballs and 17 bats 14 baseballs and 16 bats 13 baseballs and 16 bats 13 baseballs and 18 bats

OpenStudy (tkhunny):

Have you considered simply multiplying to figure the total cost of each combination? 14*20 + 17*35 = ?? Is it less than 871?

OpenStudy (anonymous):

Ohh you plug them in like 20(x)+35(x)=870 :D @tkhunny :D thank you ! I forgot

OpenStudy (tkhunny):

The substitution principle is very important. Note: It may not be EXACTLY 870. Just make sure it doesn't go over! Note: You used 'x' to mean the number of balls and the number of bats. Don't do that. 20(x) + 35(y) <= 870 is the relation to examine.

OpenStudy (anonymous):

13 baseballs and 16 bats ?

OpenStudy (tkhunny):

You tell me. Prove it. I didn't work the problem and don't intend to.

OpenStudy (anonymous):

20(13)+35(16)=820

OpenStudy (anonymous):

14 baseballs and 16bats